OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 – x2 – 20x ?

OpenStudy (nenadmatematika):

=x(x^2-x-20)=x(x-5)(x+4) so the zeros are x=0, x=5 and x=-4